OpenStudy (anonymous):

Find all the unit vectors in the plane determined by the vectors u = (-1,-2,1) and v = (2,3,0), that are perpendicular to the vector w = (-1,0,3).

OpenStudy (anonymous):

I don't know if what I did is correct, but this is what I have done: The plane determined by the vectors u and v is all the vectors in the form of au + bv So: au = a(-1,-2,1) = (-a,-2a,a) bv = b(2,3,0) = (2b,3b,0) So au + bv = (-a + 2b, -2a + 3b, a) Now, (au+bv).w = 0 (-a + 2b, -2a + 3b, a) . (-1,0,3) = 0 (-1)(-a+2b) + (0)(-2a+3b) + 3(a) = 0 a - 2b +3a = 0 4a = 2b 2a = b Sub b = 2a into (-a + 2b, -2a + 3b, a): (-a + 2b, -2a + 3b, a) =(-a + 2(2a), -2a + 3(2a), a) = (3a, 4a, a) So: a(3,4,1) is the sum of all the vectors that determine the plane Therefore, find the norm +/- sqrt(3^2 + 4^2 + 1^2) +/- sqrt(9+16+1) +/- sqrt(26) So, the unit vectors are: (3,4,1)/sqrt(26) = (3/sqrt(26), 4/sqrt(26), 1/sqrt(26)) And - (3,4,1)/sqrt(26) = (-3/sqrt(26), -4/sqrt(26), -1/sqrt(26))

OpenStudy (anonymous):

It's +/- sqrt(26) because |a|*norm = 1 to be a unit vector So +/- 1/norm is a